OpenStudy (anonymous):

Choose the equation that could be used to find two consecutive odd integers whose sum is 24

OpenStudy (anonymous):

x+x+2=24

OpenStudy (anonymous):

x+(x+2)=24

OpenStudy (anonymous):

or even x+(x-2)=24.